Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

In a certain game, you perform three tasks. You flip a

Expert replies
by swerve » Fri Nov 29, 2019 1:42 pm
In a certain game, you perform three tasks. You flip a quarter, and success would be heads. You roll a single die, and success would be a six. You pick a card from a full playing-card deck, and success would be picking a spades card. If exactly one of these three tasks is successful, then you win the game. What is the probability of winning?

A. 1/48
B. 5/16
C. 11/12
D. 11/16
E. 23/48

The OA is E

Source: Magoosh
Join the discussion
Source: — Problem Solving |

by Jay@ManhattanReview » Mon Dec 02, 2019 9:39 pm
swerve wrote:In a certain game, you perform three tasks. You flip a quarter, and success would be heads. You roll a single die, and success would be a six. You pick a card from a full playing-card deck, and success would be picking a spades card. If exactly one of these three tasks is successful, then you win the game. What is the probability of winning?

A. 1/48
B. 5/16
C. 11/12
D. 11/16
E. 23/48

The OA is E

Source: Magoosh
Say

the probability of getting a head = p = 1/2; thus, the probability of getting a tail = q = 1/2;
the probability of getting a six = x = 1/6; thus, the probability of getting any among (1, 2, 3, 4 and 5) = y = 5/6;
the probability of getting a card of spade = r = 13/52 = 1/4; the probability of getting a card other than spade = s = 39/52 = 3/4;

Given that if exactly one of these three tasks is successful, then you win the game.

The probability of winning = pys + qxs + qyr = 1/2*5/6*3/4 + 1/2*1/6*3/4 + 1/2*5/6*1/4 = 23/48

The correct answer: E

Hope this helps!

-Jay
_________________
Manhattan Review GRE Prep

Locations: GRE Classes Los Angeles | GMAT Prep Course Singapore | GRE Prep Orlando | SAT Prep Courses Toronto | and many more...

Schedule your free consultation with an experienced GMAT Prep Advisor! Click here.
Join the discussion

by Scott@TargetTestPrep » Thu Dec 05, 2019 7:02 pm
swerve wrote:In a certain game, you perform three tasks. You flip a quarter, and success would be heads. You roll a single die, and success would be a six. You pick a card from a full playing-card deck, and success would be picking a spades card. If exactly one of these three tasks is successful, then you win the game. What is the probability of winning?

A. 1/48
B. 5/16
C. 11/12
D. 11/16
E. 23/48

The OA is E

Source: Magoosh
We can let A = event of getting heads when flipping the quarter, B = event of getting a six when rolling the die and C = event of getting a spades card, and use the following formula:

P(A or B or C) = P(A) + P(B) + P(C) - P(A and B) - P(A and C) - P(B and C) + P(A and B and C)

P(A or B or C) = 1/2 + 1/6 + 1/4 - (1/2 x 1/6) - (1/2 x 1/4) - (1/4 x 1/6) + (1/2 x 1/6 x 1/4)

P(A or B or C) = 11/12 - 1/12 - 1/8 - 1/24 + 1/48

P(A or B or C) = 11/16

Alternate Solution:

We notice that P(success) + P(failure) = 1; therefore, P(success) = 1 - P(failure). Let’s find P(failure).

The only way we fail in this game is if we get tails from the quarter flip AND not get a six from the die roll AND not get a spade from the card draw. Therefore,

P(failure) = 1/2 x 5/6 x 3/4 = 15/48 = 5/16

Thus, P(success) = 1 - P(failure) = 1 - 5/16 = 11/16

Answer: D

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion

by [email protected] » Fri Dec 06, 2019 11:17 am
Hi All,

We're told that in a certain game, you perform three tasks. You flip a quarter, and success would be heads. You roll a single die, and success would be a six. You pick a card from a full playing-card deck, and success would be picking a spades card. We're told that if you perform EXACTLY ONE of these three tasks is successful, then you win the game (meaning that any other outcomes - re: 0, 2 or 3 successes would NOT be a "win"). We're asked for the probability of winning the game.

Since there are 2 outcomes with a coin, 1/2 of the outcomes are successes and 1/2 are not
Since there are 6 outcomes with rolling a die, 1/6 of the outcomes are successes and 5/6 are not
Since there are 52 outcomes with choosing a card (and 13 cards are spades), 1/4 of the outcomes are successes and 3/4 are not

There are 3 ways to 'win' the game:
1) Win the coin, Lose the die, Lose the card = (1/2)(5/6)(3/4) = 15/48
2) Lose the coin, Win the die, Lose the card = (1/2)(1/6)(3/4) = 3/48
3) Lose the coin, Lose the die, Win the card = (1/2)(5/6)(1/4) = 5/48

Total probability of winning the game = 15/48 + 3/48 + 5/48 = 23/48

Final Answer: [spoiler=]E[/spoiler]

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion